OpenStudy (anonymous):

what is 49/50 as a percentage

OpenStudy (anonymous):

98%

OpenStudy (anonymous):

if you can't use a calculator then make the denominator 100 for easy math if you don't know it off the top of your head.

OpenStudy (anonymous):

Okay, so percent is out of 100 right? How would you convert 49/50 to something/100?

OpenStudy (anonymous):

@adnane.benali Please help the asker find the answer, but don't answer for them. That way they can best learn how to figure it out on their own.

OpenStudy (anonymous):

Apologies I gave away the answer. yes, as the other said, 49 out of 50 is (49x2) out of (50x2) ==> 98 out of 100.

OpenStudy (anonymous):

Another way I usually do it if it's complicated numbers is this: I write down the values given, in this case: 49 ................ 50 then underneath it I write down a question mark, what I need to calculate, and 100 as follows: 49................50 ?...................100 then I cross multiply 49 by 100 and divide by 50. Imagine a cross, in order to solve for the "?" so ? = 49 x 100 (the diagonal line from 49 to 100) divided by 50 (imagine a diagonal line from 50 to the ?. ? = 4900/50 = 98 try this with complicated numbers like... express in % the following 34/76 so I do: 34...................76 ?......................100 then: ? = 34x100/76 = 3400/76 = 44.7% Hope this helps.